OpenStudy (anonymous):

Find the altitude of an equilateral triangle, if each side of the triangle has a length of 20 meters. a) 10 squareroot 3 b) 10 squareroot 2 c) 10 meters d) 20 meters

OpenStudy (anonymous):

pythagorean's theorem. a^2 +b^2 = c^2

OpenStudy (anonymous):

a) 10 squareroot 3

OpenStudy (anonymous):

not pythagorean

OpenStudy (anonymous):

Draw it out. You want the altitude. Split the triangle in half.

OpenStudy (anonymous):

you find that the hypotenuse is 20 and the base 10.

OpenStudy (anonymous):

plug it in. c is the hypotenuse. the base can be either a or b. then solve for the variable.

OpenStudy (anonymous):

step 1. a^2 +b^2 = c^2 step 2. 10^2 + b^2 = 20^2 step 3. b^2 = 20^2 - 10^2 step 4. b^2 = 400 - 100 step 5. b^2 = 300 step 6. b = sqrt (300) step 7. b = 10*sqrt(3) [Here I pulled out 100 from 300 and rooted it, giving me 10]

OpenStudy (anonymous):

@moneybird yes it is pythagorean's theorem. I just proved it

OpenStudy (anonymous):

@moneybird giving answers doesn't help anybody...You must provide steps. Otherwise they won't be able to answer the question properly

OpenStudy (anonymous):

Well i don't have any steps i guessed the answer in my head

OpenStudy (anonymous):

Lol guessing. Horrible :[ .

OpenStudy (anonymous):

Guessing is good when you can guess the right answer

OpenStudy (anonymous):

You should learn the proper steps though :[